OpenStudy (anonymous):

Evaluate the following expressions when c= -7 and d= 4.

OpenStudy (anonymous):

\[\sqrt{d}-dc ^{2}\]

hartnn (hartnn):

so when u put c=-7 and d= 4 u get \(\sqrt4-4(-7)^2\) =2-4(49) can u solve further ?

OpenStudy (anonymous):

yes thank you

hartnn (hartnn):

welcome :)